OpenStudy (anonymous):

Solve 1= a(10-5)^2 + 3

OpenStudy (anonymous):

(quadratic Equation)

OpenStudy (ash2326):

Is this the question? \[1=a(10-5)^2+3\]

OpenStudy (lgbasallote):

10-5 is 5 5^2 is 25 so it becomes 1 = 25 a + 3 -2 = 25a a = -2/25

OpenStudy (lgbasallote):

is that right question??

OpenStudy (lgbasallote):

coz i dont see quadratic there